The area under restoring force graph gives what __________.

Grade:11

1 Answers

Vikas TU
7 years ago
Consider that restoring force F = kx.
x as abscissa and F is at ordinate.
F = ma 
Let at particular some x and y coordinates we consider small area that is:
dA = a*dx
     = vdv/dx dx
      =v dv
Integerating both sides, we get
A = (1/2)v^2
Now this is the area under acceleration and displacement x.
For force multiplying with m for F = ma we get (1/2)mv^2
that is K.E.
Hence, 
 
The area under restoring force graph gives Kinetic Energy.